for copper the variation with temperature for heat capacity at constant pressure is Cp=22.46+(00.628) calculate the change of the molar enthalopy when temperature change from293 to 393


1
Expert's answer
2021-12-22T14:16:05-0500

Cp = dH/dT at constant pressure

Cp = ∆H/(T2 - T1)

∆H = Cp x (T2 - T1) = (22.46+(00.628)) x (393 - 293) = 2308.8 J